OpenStudy (anonymous):

Use the division rule to simplify the following: 3x^4/y^3 divide 6x^2/y^2

OpenStudy (anonymous):

a/b divided by c/d=ad/bc

OpenStudy (anonymous):

\[\left(\frac{3 x^4}{y^3}\right)/\left(\frac{6 x^2}{y^2}\right)=\frac{3 x^4}{y^3}*\frac{y^2}{6 x^2}=\frac{x^2}{2 y} \]